OpenStudy (anonymous):

Jim halves the distance between himself and a sound source. What is the change in decibels of the sound he hears?

OpenStudy (anonymous):

db = 10log(intensity of sound/threshold of human hearing) Intensity = Power/Area Intensity = Power/4 pie r^2 Power = Intensity x 4 x pie x r^2 equating power Intensity(originally) x 4 x pie x r^2 = Intensity(at a distance 2r) x 4 x pie x (2r)^2 2r because he moves twice as far away divide 4 x pie lo intensity originaly, la intensity after moving Io x r^2 = Ia x (2r)^2 Io = Ia x 4 lets say he was originally 1m, and moves to 2m change in dB = dbo - dba = 10log(1/10^-12) - 10log(1/4 x 10^-12) = 10log(4) =6 db change

OpenStudy (anonymous):

ive just realised i did it as if he starts close then moves twice as far away. the answer is exactly the same for if he started far away and then moves close.